OpenStudy (anonymous):

@jessiehuang_ Which equation could be used to solve this problem? The sum of two consecutive integers is 141. Find the integers. A. n + 1= 141 B. 2n + 2 = 141 C. 2n = 141 D. 2n + 1 = 141

OpenStudy (anonymous):

@jessiehuang_

OpenStudy (anonymous):

always tag @jessiehuang_

OpenStudy (anonymous):

i don't remember this. i'd go with d.

OpenStudy (anonymous):

do u want me to swich than

OpenStudy (anonymous):

i'm pretty sure it's d

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

your

OpenStudy (anonymous):

Which equation could be used to solve this problem? The sum of two consecutive odd integers is 124. Find the integers. A. n + 3 = 124 B. 2n + 4 = 124 C. 2n + 2 = 124 D. 2n + 1 = 124

OpenStudy (nincompoop):

an example of two consecutive numbers are 2, 3 4, 5 5, 6 from the pattern, you will notice that whatever number you started (it maybe represented by n or x) you just have to add 1 x+1 or n + 1

OpenStudy (anonymous):

Which equation could be used to solve this problem? The sum of two consecutive odd integers is 124. Find the integers. A. n + 3 = 124 B. 2n + 4 = 124 C. 2n + 2 = 124 D. 2n + 1 = 124

OpenStudy (nincompoop):

post your entire homework here so we can do it for you

OpenStudy (anonymous):

i'm sorry i can't remember how to do consecutive integers. forgive me @billguy

OpenStudy (anonymous):

its cool ill go to another 1

OpenStudy (nincompoop):

it says ODD integers

OpenStudy (anonymous):

im going to another close and reopen

OpenStudy (alekos):

Consecutive integers "D" Consecutive odd integers "none of the above"